### v1

# The Cognitive Foundations of Economic Exchange: A Modular Framework Grounded in Behavioral Evidence

Abstract
A key challenge in multi-agent AI is modeling social cooperation under realistic behavioral constraints. Many foundational concepts in economics and ethics—such as ”trust” or ”morality”—are often defined informally, without operational criteria or cognitive grounding, which limits their testability and implementation in artificial agents. Drawing on converging empirical evidence from primate behavior, infant cognition, and economic anthropology, we propose a conceptual framework composed of three cognitively minimal mechanisms: individual recognition, reciprocal credence, and cost–return sensitivity. This framework reframes trust as a graded cognitive expectation, providing a simulateable basis for reciprocal exchange in artificial agents, and enabling the bottom-up emergence of scalable cooperation and institutional dynamics.
Machine Learning, ICML
Figure 1: Three Core Cognitive Mechanisms—Individual Recognition, Reciprocal Credence, and Cost–Return Sensitivity—as Behavioral Primitives for Simulating Reciprocal Exchange in Artificial Agents.
1 Introduction
A persistent gap in economic simulation is the lack of behavioral grounding: most models begin with institutions—markets, contracts, payoff matrices—while assuming the cognitive and social mechanisms that make such systems viable. This limits both the explanatory power and simulateability of artificial economic agents.
This issue is mirrored in broader AI+Society research. Concepts like “trust,” “value,” and “cooperation” are often treated as abstract labels, operationalized only at the outcome level. Without grounding these constructs in observable mechanisms, models risk simulating the surface of economic life without modeling its origins.
In this paper, we argue that the foundations of economic exchange—and by extension, scalable cooperation—can be traced to three cognitively minimal, biologically grounded mechanisms: individual recognition, reciprocal credence, and cost–return sensitivity. These capacities are observed across human and nonhuman primates, and support sustained reciprocal exchange without requiring symbolic trust or formal enforcement.
Anthropological accounts have long shown that early human exchange was not rooted in barter or equivalence, but in socially embedded reciprocity. Systems like the Kula ring (Malinowski, 2013) emphasized obligation and alliance over transaction. Similarly, primatology has documented robust reciprocal behaviors among chimpanzees and bonobos—grooming (De Waal, 1997), food sharing, and coalition—indicating that the behavioral foundations of reciprocal exchange are observable across species and precede formal market systems.
We propose a modular framework built from these three mechanisms, designed to serve as behavioral primitives in economic simulation. This approach not only bridges cognitive science and institutional theory, but also enables bottom-up modeling of how debt, credit, and monetary behavior may emerge from repeated social interaction.
Our contribution.
We offer a theoretical realignment: a cognitively grounded account of reciprocal exchange that clarifies how economic systems can emerge from social behavior rather than institutional design. Specifically, we:
  • Identify three minimal mechanisms—individual recognition, reciprocal credence, and cost–return sensitivity—as simulateable behavioral primitives for modeling the emergence of economic exchange;
  • Synthesize evidence from primate behavior, infant cognition, and economic anthropology to support these mechanisms across species and developmental stages;
  • Reframe “trust” as a scalar, simulateable expectation—reciprocal credence—rather than a moral abstraction;
  • Reinterpret key findings from behavioral economics—such as inequity aversion, reciprocal cooperation, and altruistic punishment—as consistent expressions of underlying biological mechanisms, rather than anomalies to rational choice theory.
This work reframes economic modeling not as a top-down implementation of rules, but as a bottom-up emergence of institutions from cognitive interaction.
Ethical Statement
Importantly, this work does not rely on evolutionary explanations. While we draw on behavioral evidence from primates and human infants, our goal is not to claim innate or adaptive origins of exchange. Rather, these cases serve as empirical constraints to identify minimal cognitive mechanisms sufficient for reciprocal behavior. Our account is grounded in behavioral plausibility, not evolutionary teleology.
2 Related Work
2.1 Agent-Based Social Simulation
Recent advances in multi-agent language models have produced systems capable of planning, negotiation, and scripted cooperation (Park et al., 2023; Li et al., 2023). However, these models typically lack behavioral grounding: they simulate interaction as token-level dialogue without modeling the internal mechanisms that make social exchange stable—such as partner memory, reciprocity tracking, or cost–benefit sensitivity.
While some frameworks include persistent memory modules or heuristic-based reasoning, these components often lack principled constraints. Agents either memorize indiscriminately or update state in ways that are unstable or opaque, making it difficult to simulate long-term reciprocal dynamics. As a result, apparent cooperation often emerges from prompt bias or hardcoded behavior, not from simulateable social inference.
This gap has two key consequences: (1) It limits the explanatory value of multi-agent simulations as models of human social behavior, and (2) it prevents systematic exploration of how early exchange mechanisms scale into institutional forms. Our framework addresses this by identifying minimal, biologically plausible mechanisms that can be embedded in LLM agents to support grounded, dynamic reciprocity over time.
2.2 Reciprocal Behavior in Nonhuman Primates
Research in primatology provides critical insight into the behavioral foundations of reciprocity. Species such as chimpanzees and bonobos engage in reciprocal behaviors across contexts like food sharing, grooming, and coalition support. For instance, de Waal (De Waal, 1997) illustrated how chimpanzees form long-term social alliances maintained through reciprocal exchange, and Brosnan et al. (Brosnan & De Waal, 2003) demonstrated that primates are sensitive to fairness and inequity in outcomes.
While these studies have richly documented the behavioral forms of reciprocity, they often stop short of linking such behaviors to the broader architecture of economic interaction. The question of how these mechanisms scale or evolve into structured systems of exchange remains relatively underexplored.
Our framework builds on these findings by situating primate reciprocity within a cognitive-behavioral model. We argue that capacities such as individual recognition, memory of interaction history, and fairness sensitivity form a biologically grounded substrate upon which more complex human economic systems could emerge.
2.3 Social Exchange Without Markets
Anthropological accounts have long challenged the barter-origin myth that underpins classical economic theory. Foundational works such as The Gift (Mauss, 2024) and Stone Age Economics (Sahlins, 2013) argue that early exchange was not driven by market logic, but embedded in webs of social relationships, mutual obligation, and symbolic prestige.
These accounts have been instrumental in shifting attention away from transactional equivalence and toward socially embedded forms of value. Yet, they often bracket the question of what kinds of cognitive and behavioral mechanisms make such systems feasible. In other words: what allows individuals to engage in sustained, deferred, and socially contingent exchange?
Our framework complements these anthropological insights by identifying the minimal behavioral architecture that enables reciprocity to stabilize across time and social partners. We argue that reciprocity is not merely a cultural invention, but a behaviorally tractable structure observable across species.
2.4 From Behavioral Anomalies to Cognitive Substrates
Behavioral economics has significantly advanced our understanding of social preferences, particularly in contexts of fairness, cooperation, and reciprocity. Paradigms such as the Ultimatum Game (Güth et al., 1982), Trust Game (Berg et al., 1995), and Public Goods Game (Fehr & Gächter, 2002) reveal systematic deviations from self-interested rationality, with participants rejecting unfair offers, rewarding cooperation, and punishing free-riders.
These empirical patterns have led to influential models of “inequity aversion” (Fehr & Schmidt, 1999) and “strong reciprocity” (Gintis, 2000), which posit internalized fairness norms and preferences for norm enforcement. However, these models typically treat the relevant capacities—such as fairness sensitivity or social tracking—as given, rather than as targets of explanation.
In contrast, our framework treats these preferences as emergent outcomes of specific cognitive-social mechanisms: individual recognition, memory of interaction, and cost–return evaluation. By shifting the focus from outcomes to underlying capacities, we aim to bridge behavioral economic observations with a biologically grounded account of exchange behavior.
Figure 2: Evidence from anthropology, primatology, and developmental psychology suggests that the behavioral foundations of economic exchange lie in reciprocity—not barter or market logic.
3 Background: Behavioral Origins of Exchange
3.1 The Myth of Barter and the Reciprocal Foundations of Exchange
The textbook account of economic origins often begins with barter: the notion that early humans exchanged goods directly in the absence of currency, with markets and money arising later to reduce transactional friction. This narrative, while intuitively appealing, has little support in either ethnographic or historical records.
As Sahlins argued in Stone Age Economics (Sahlins, 2013), early exchange systems were not structured around equivalence, but around relational modes of reciprocity—generalized, balanced, and negative. These forms differ not in rationality but in social distance and temporal orientation: generalized reciprocity, typical among kin, involved open-ended giving without expectation of immediate return; balanced reciprocity, among peers, entailed roughly equivalent exchanges over time; and negative reciprocity, often among rivals or strangers, reflected opportunistic or exploitative behavior.
A paradigmatic ethnographic example is the Kula ring, described by Malinowski (Malinowski, 2013), in which ceremonial armbands and necklaces circulate across Trobriand Islands through long-term, prestige-based exchanges. These objects held little utilitarian value, yet were exchanged with strict directional patterns, social memory, and partner exclusivity. What appears as a symbolic ritual is, upon closer inspection, a complex system of relational tracking, reputational assessment, and obligation extension—precisely the kinds of cognitive demands that structure early forms of exchange.
Mauss’s The Gift (Mauss, 2024) further formalized this insight by showing how gift-giving in archaic societies is embedded in a triadic obligation: to give, to receive, and to reciprocate. The gift is not merely a transfer of value, but a means of sustaining long-term social bonds. When viewed from this angle, early exchange is not the product of impersonal rational choice, but of biologically and socially grounded behaviors—tracking individuals, remembering interactions, maintaining alliances, and adjusting for asymmetrical returns.
These foundations suggest that exchange predates markets not just historically, but cognitively. What emerges from these systems is not merely institutional scaffolding, but a structured behavioral logic—one that we formalize in the following section.
3.2 The Three Core Mechanisms of Biological Reciprocity
If the earliest forms of economic interaction were not driven by impersonal barter, then what behavioral dynamics made sustained exchange possible? We argue that reciprocal exchange emerged not from a need to equalize value, but from the ability to maintain accountability and continuity across repeated social interactions.
These early exchanges were not market transactions, but embedded social patterns shaped by memory, mutual responsiveness, and partner-specific interaction history. Exchange, in this view, arises not from price formation, but from behavioral infrastructure: the cognitive and social capacities that allow individuals to track others, assess cooperative value, and anticipate reciprocity.
To formalize these behavioral foundations, we propose a cognitively grounded framework composed of three simulateable mechanisms:
  • Individual recognition: enabling agents to identify and re-engage specific social partners over time;
  • Reciprocal credence: a scalar, updateable expectation that cooperation will be returned;
  • Cost–return sensitivity: allowing agents to regulate cooperative behavior based on dynamic payoff asymmetries.
These three capacities are supported by converging evidence from nonhuman primates and human infants, suggesting they are both biologically plausible and developmentally early. Importantly, they are not merely descriptive—they are simulateable as modular components in multi-agent systems.
While primate studies have richly documented reciprocal behaviors—grooming, food sharing, coalition support—the cognitive prerequisites for sustaining such behavior are often left implicit. As (Schino & Aureli, 2010) note, reciprocity is frequently treated as a local response pattern, rather than a structured process involving memory and expectation. Yet reciprocal exchange, by definition, requires continuity. Without recognition or anticipation of future interaction, trust-like behavior cannot functionally emerge.
By defining these mechanisms as behavioral primitives, we offer a bottom-up model of economic emergence—one that grounds institutional systems not in rules, but in biologically plausible, cognitively minimal substrates of social interaction.
4 Theoretical Framework: Simulateable Primitives for Economic Exchange
Existing efforts in agent-based cooperation often rely on engineered incentives, fixed strategy spaces, or symbolic approximations of concepts like “trust.” What remains missing is a biologically grounded, simulateable account of the behavioral capacities that make reciprocal exchange possible.
We propose three minimal mechanisms—individual recognition, reciprocal credence, and cost–return sensitivity—as a compositional framework for simulating the emergence of economic exchange. These mechanisms are:
  • Cognitively minimal: each is supported by empirical data from nonhuman primates and human infants;
  • Behaviorally tractable: each has observable correlates in real-world cooperation;
  • Simulateable: each can be implemented in agent-based systems as a decision module or update function.
This framework allows social exchange to be modeled from the bottom up—without relying on predefined markets, institutions, or payoff matrices.
4.1 Individual Recognition
We define individual recognition as the ability to identify and distinguish specific agents across time, enabling the tracking of prior social interactions. This capacity underlies the first and most basic source of reciprocal credence: memory of positive experiences with known others. Without recognition, no information about past cooperation can be meaningfully retained or used to guide future behavior.
Among nonhuman primates, this capacity is especially robust. A recent study shows that chimpanzees and bonobos remember familiar individuals after decades of separation, and preferentially attend to those with whom they had positive social histories (Lewis et al., 2023). These findings show that individual recognition in apes is long-term and identity-specific, enabling the tracking of social partners across time.
In nonhuman primates, individual recognition enables partner-specific cooperation. In rope-pulling tasks, chimpanzees adjusted their actions based on the identity and prior behavior of their partners—pausing, glancing, or gesturing to initiate coordination (Hirata & Fuwa, 2007). Dyads with higher food-sharing tolerance were also more successful in joint tasks, indicating that cooperative memory is structured around known social partners(Melis et al., 2006b).
In humans, this ability emerges early. Infants as young as 14 months selectively help or share with individuals who have previously acted prosocially (Dunfield et al., 2011). Such findings suggest that recognition, even in its early form, is tied to social history and used to guide future cooperative behavior.
4.2 Reciprocal Credence
The term “trust” is widely used across disciplines, yet rarely defined with precision. In both everyday language and scientific literature, it is invoked to describe moral commitment, emotional closeness, institutional reliability, and behavioral expectation—often interchangeably. This conceptual ambiguity makes trust difficult to operationalize, especially in comparative, developmental, or artificial settings.
To address this gap, we introduce the concept of reciprocal credence: a cognitively grounded, graded expectation that another agent will respond to cooperation with cooperation. Reciprocal credence does not assume moral virtue or social obligation. Instead, it arises from informational cues—typically based on one of two elements: (1) past prosocial behavior directed toward the self or others, or (2) contextual signals that make cooperative return likely. It is not a feeling, but a function: a decision-weighting variable that governs the willingness to enter or continue reciprocal exchange.
Unlike categorical trust, reciprocal credence is inherently scalar: it reflects how likely an agent believes cooperation will be reciprocated, not whether it will be or not. Treating reciprocal credence as a scalar variable is essential to modeling cooperation as a dynamic, context-sensitive behavior.
We distinguish three primary sources of reciprocal credence:
  1. 1.
    Direct positive interaction history, where an individual has experienced prior cooperative behavior, prosocial engagement, or emotional bonding with another.
  2. 2.
    Reputational inference, where the individual observes or learns about others’ cooperative behavior.
  3. 3.
    Role-based expectation, where social roles or institutional norms lead to default expectations of reciprocity.
Each source varies in reliability, stability, and cognitive cost, but all contribute to the emergence of trust-like behavior without requiring moral commitment. Together, these sources provide the informational basis for reciprocal credence—a flexible, graded expectation that others will act in ways conducive to cooperation.
Empirical studies support the differential emergence of these sources across species and developmental stages. In chimpanzees, reciprocal cooperation is consistently observed when individuals have direct histories of prosocial interaction, with little evidence for reputation-based inference or role-based expectation. Similarly, human infants display early sensitivity to agents who have helped them directly, yet show limited use of third-party observation or role generalization. These patterns suggest that direct positive interaction is the foundational layer of reciprocal credence, while the other two sources emerge developmentally and culturally.
Across species, direct prosocial interaction serves as the most robust and foundational basis for reciprocal credence. In wild chimpanzees, grooming is exchanged not merely in-the-moment, but over extended time horizons. Longitudinal data reveal that grooming given and received is significantly correlated across dyads, even with delays spanning several days—indicating that chimpanzees maintain memory of past interactions and calibrate future behavior accordingly (Gomes et al., 2009). Experimental studies further show that chimpanzees prefer to collaborate with previous cooperative partners, suggesting they encode individual-specific social history and deploy it strategically in cooperative contexts (Melis et al., 2006a).
Human infants exhibit a comparable early capacity. By three years of age, children are less likely to help individuals who have previously acted with harmful intent—even when the harm was directed at a third party, and even if no harm was successfully done. This selective avoidance indicates that infants track not only outcomes but also the moral valence of agents’ prior actions when deciding to engage prosocially (Vaish et al., 2010).
Together, these findings support the primacy of direct positive interaction in generating reciprocal credence. Unlike inference or abstraction, this source relies on lived experience—anchoring cooperative expectations in observable history, not hypothetical roles or indirect cues.
A note on “trust”.
It is also important to distinguish social trust from what might be better described as functional dependency. Contemporary systems such as ChatGPT or Google are often described as being “trusted,” when in reality they are simply relied upon due to their consistent output, broad adoption, and lack of failure. This perceived “trust” is better understood as a product of usability, network effects, and repeated exposure — not reciprocal social accountability.
A similar semantic conflation is common in financial discourse, where “trust in money” or “trust in banks” is frequently cited as foundational to economic stability. However, what is usually meant in such contexts is not interpersonal trust, but confidence in the continuity of institutional enforcement, liquidity guarantees, or the absence of systemic failure.
4.3 Cost–return sensitivity
We define cost–return sensitivity as the capacity to evaluate whether engaging in cooperative behavior yields a net benefit over time. Unlike categorical heuristics (e.g., always cooperate or always defect), this sensitivity enables agents to condition their participation on the payoff structure of ongoing interactions. It allows withdrawal from exploitative relationships and reinforcement of beneficial ones, serving as a regulatory mechanism for reciprocal dynamics.
This capacity need not involve explicit calculation; approximate assessments—such as tracking recent imbalances in giving and receiving—are sufficient to support adaptive behavior. For example, chimpanzees are less likely to groom or share with partners who have failed to reciprocate in the past, even if those costs are not immediate or symmetrical. Similarly, infants gradually reduce helping behavior toward agents who consistently fail to reciprocate attention or aid.
Such sensitivity ensures that reciprocal exchange remains robust under uncertainty, preventing exploitation while maintaining flexibility in dynamic social environments.
In nonhuman primates, cost–return sensitivity manifests in both prosocial and retaliatory contexts. Chimpanzees engage in grooming and service-like exchanges where helping behavior depends on previous benefits received from others (De Waal, 1997). They are also capable of negative reciprocity: retaliating against individuals who have stolen from them, even when such retaliation yields no direct benefit—suggesting an expectation of equitable return and a sensitivity to intentional harm (Jensen et al., 2007).
In humans, early traces of this mechanism appear in infancy. By 18 months, children selectively direct helping or sharing behavior toward agents who previously acted fairly or cooperatively (House et al., 2013). These decisions are not based on abstract principles of justice, but on perceived patterns of cost and return—who gave, who withheld, and under what circumstances.
Taken together, these findings suggest that cost–return sensitivity enables organisms to calibrate cooperative investment, maintaining reciprocity in the absence of formal institutions or explicit contracts. As a building block of economic cognition, it governs the micro-dynamics of exchange long before the emergence of price or property.
5 Possible Implementation via Memory Design
While each proposed mechanism is tractable in isolation, integrating them into a unified architecture remains an open challenge—particularly due to current limitations in memory design for LLM-based agents. Most existing multi-agent frameworks lack persistent, structured memory, resulting in shallow social modeling and limited capacity for behavioral adaptation.
We propose that each behavioral primitive—individual recognition, reciprocal credence, and cost–return sensitivity—can be realized through targeted extensions to agent memory and retrieval mechanisms:
  • Individual recognition can be implemented using identity-specific memory slots or embedding-indexed retrieval systems, enabling agents to store and retrieve interaction histories with named partners. This supports continuity in agent behavior and the ability to condition future responses on partner identity.
  • Reciprocal credence may be represented as an internal latent variable or scalar annotation associated with each agent in memory (e.g., “Agent X: credence = 0.7”), and updated through lightweight heuristics (e.g., recency-weighted prosocial acts). This allows LLM agents to infer, represent, and update partner expectations without requiring explicit reward functions.
  • Cost–return sensitivity can be implemented through structured memory logs of past cooperative attempts and outcomes. Agents may track cumulative give–take asymmetries and adjust future investments accordingly. Prompts can surface this information as part of decision rationale (e.g., “Agent Y has taken without giving for three turns”).
These strategies enable biologically inspired reciprocity to be modeled using memory-centric extensions, rather than reinforcement learning or symbolic planning. By shifting the burden of social behavior onto structured memory design and prompt-level reasoning, this framework opens a tractable path toward more socially competent LLM agents.
6 Implications and Discussion
6.1 Limitations of the Framework
Comparative Methodology.
Comparative psychology and developmental psychology use fundamentally different experimental paradigms. While primate studies often capture interactive, ecologically valid behaviors, human infant research typically relies on simplified, highly constrained designs. This asymmetry limits the granularity with which we can compare capacities across species.
Institutional Complexity.
Although we argue that modern exchange institutions build upon a biologically grounded behavioral substrate, we do not claim that such mechanisms alone are sufficient to explain the emergence of formal systems like currency, interest-bearing debt, or taxation. These likely require additional cultural, historical, and symbolic developments.
Social Scale and Enforcement.
Our framework focuses on the dyadic and small-group foundations of reciprocal exchange, where individuals recognize and respond to the behavior of specific partners. It does not yet account for mechanisms that support large-scale, anonymous cooperation—such as third-party punishment, norm internalization, or reputation management across broader social networks. These are critical features of modern institutions, and extending the framework to address how such systems emerge from or layer upon basic social cognition remains a key direction for future work.
6.2 Theoretical Implications
From Institutions to Cognition.
Our work reorients the study of exchange from an institutionalist or symbolic perspective toward a cognitive-behavioral foundation. Rather than treating markets, money, or debt as cultural inventions that enabled exchange, we argue that exchange itself is grounded in prior cognitive and social capacities—capacities that predate and scaffold institutionalization.
Rethinking Trust as System Robustness.
The term “trust” is often invoked in discussions of economic exchange, yet it bundles together distinct ideas—from interpersonal reliability to confidence in systems. We argue that this vagueness limits its explanatory value.
To address this, we introduce reciprocal credence: a graded, biologically grounded expectation that cooperation will be returned. It emerges from recognition, interaction history, and cost–return sensitivity—not from moral obligation. In contrast, large-scale “trust” in systems like money or platforms is better understood as confidence in robustness, not social accountability. This reframing clarifies how reciprocal behavior scales without assuming institutional trust.
Cross-Species Foundations of Exchange.
By aligning primate cooperation studies with human developmental data and ethnographic accounts of non-monetary exchange, we provide a cross-species bridge for modeling the origins of economic behavior. This integrative view challenges the idea that complex exchange systems are uniquely human, and instead locates their roots in broader social cognition.
Unifying Behavioral Economics through Cognitive Foundations.
Many findings in behavioral economics—such as fairness preferences, reciprocal cooperation, and sensitivity to framing or loss—are often treated as deviations from rational choice theory. Yet evidence from primatology and developmental psychology suggests that these behaviors reflect consistent biological patterns: they emerge early in human development, recur across great apes, and follow predictable social-cognitive constraints. Rather than anomalies or patchwork heuristics, these tendencies may be better understood as structured expressions of an underlying behavioral logic—one shaped by the cognitive demands of navigating social exchange.
6.3 Future Work
6.3.1 Operationalizing Economic Emergence in Agent Simulations
A key direction for future research is to instantiate the three proposed behavioral primitives—individual recognition, reciprocal credence, and cost–return sensitivity—as explicit, modular components in LLM-based multi-agent systems.
This involves not only defining each mechanism’s internal representation (e.g., persistent memory slots for identity recognition, scalar credence values for reciprocity tracking, and interaction-weighted cost–return summaries), but also studying how they interact dynamically across time and partners.
We propose that future simulations systematically test:
  • How memory constraints (e.g., limited identity tracking) affect the emergence of sustained cooperation;
  • Whether agents with scalar credence variables spontaneously differentiate partners and regulate prosocial behavior;
  • Under what conditions cost–return asymmetries trigger withdrawal, retaliation, or symbolic substitutes (e.g., tokens, credit markers).
Such experiments can help identify the minimal memory and inference structure necessary to support scalable exchange. More broadly, this line of work offers a concrete path toward simulating institutional emergence—from micro-level reciprocity to meso-level norms and proto-economic artifacts—grounded in biologically plausible cognition.
6.3.2 Simulating Institution Formation via Behavioral Primitives
A promising direction for future work is to simulate the emergence of economic institutions—such as debt, symbolic exchange, and trust networks—by embedding behavioral primitives into artificial agents. These primitives, derived from comparative psychology and developmental research, include individual recognition, reciprocal credence, and cost–return sensitivity—as the minimal substrate for modeling scalable cooperation. Memory and bonding may be embedded within the learning processes that govern credence formation.
With the rise of large language model (LLM)-based agents and embodied multi-agent simulators (e.g., Habitat, ThreeDWorld, AutoGen), it is now possible to explore whether agents endowed with these primitives can spontaneously develop systems of reciprocal exchange, role specialization, symbolic debt, or proto-monetary behaviors. Such experiments would help identify the minimal substrate for scalable cooperation and provide a generative framework for building socially competent artificial agents.
Importantly, this approach also enables us to test under what conditions behavioral reciprocity transitions into institutionalization: for example, whether limited memory capacity leads to symbolic tokens, or whether increasing agent density necessitates generalized trust representations. Simulation environments offer a new way to bridge cognitive models of exchange with the design of future artificial social systems.
6.3.3 Exchange among unfamiliar agents
One unresolved but critical question concerns the possibility of reciprocal exchange among unfamiliar agents. While most existing studies—whether in primatology, developmental psychology, or economic anthropology—focus on cooperation within socially familiar contexts, little is known about how exchange unfolds in the absence of recognition, shared norms, or interaction history. Transactions between unfamiliar individuals, especially without institutional enforcement, may require distinct cognitive or communicative scaffolding that is not yet well understood. We highlight this not as a limitation of the current framework, but as an open frontier that future research must address.
7 Conclusion
This paper proposes a cognitively grounded framework for modeling the behavioral origins of economic exchange in multi-agent systems. Instead of relying on institutional assumptions or symbolic abstractions, we identify three simulateable primitives—individual recognition, reciprocal credence, and cost–return sensitivity—that support sustained reciprocal cooperation in the absence of formal mechanisms.
Our framework addresses a central modeling challenge: how to operationalize constructs such as “trust” or “morality,” which are foundational in human societies but often lack consensus definitions or concrete implementation paths in computational systems. Drawing on evidence from primate behavior, infant cognition, and economic anthropology, we reframe these abstract concepts in terms of observable, cognitively minimal mechanisms—providing a tractable basis for implementation in artificial agents.
This perspective encourages a shift from top-down economic modeling to bottom-up simulation, where complex structures emerge from agent interaction rather than being imposed. Future work can embed these primitives into large-scale multi-agent environments to explore the spontaneous formation of scalable cooperation and exchange systems.
Impact Statement
This paper proposes a cognitively grounded framework for simulating the behavioral origins of economic exchange, with potential applications in multi-agent modeling, artificial social systems, and foundational research in AI + Society. While the work is theoretical and does not involve data collection or deployment, it provides biologically plausible primitives that could improve the interpretability and social competence of agent-based systems. We foresee no foreseeable negative societal consequences or ethical risks associated with this research.

### v2

@@
 Drawing on converging empirical evidence from primate behavior, infant cognition, and economic anthropology, we propose a conceptual framework composed of three cognitively minimal mechanisms: individual recognition, reciprocal credence, and cost–return sensitivity.
-This framework reframes trust as a graded cognitive expectation, providing a simulateable basis for reciprocal exchange in artificial agents, and enabling the bottom-up emergence of scalable cooperation and institutional dynamics.
+This framework reframes trust as a graded cognitive expectation, providing a simulateable basis for reciprocal exchange in artificial agents, enabling the bottom-up emergence of scalable cooperation, proto-economic norms, and institutional structure.
 Machine Learning, ICML
@@
 Without grounding these constructs in observable mechanisms, models risk simulating the surface of economic life without modeling its origins.
-In this paper, we argue that the foundations of economic exchange—and by extension, scalable cooperation—can be traced to three cognitively minimal, biologically grounded mechanisms: individual recognition, reciprocal credence, and cost–return sensitivity.
+In this paper, we argue that the foundations of economic exchange—and by extension, scalable cooperation—can be traced to three cognitively minimal, behaviorally grounded mechanisms: individual recognition, reciprocal credence, and cost–return sensitivity.
 These capacities are observed across human and nonhuman primates, and support sustained reciprocal exchange without requiring symbolic trust or formal enforcement.
@@
 We offer a theoretical realignment: a cognitively grounded account of reciprocal exchange that clarifies how economic systems can emerge from social behavior rather than institutional design.
-Specifically, we:
+We argue that reciprocity is not just a component of cooperation, but the minimal cognitive substrate from which stable social structures emerge.
   • Identify three minimal mechanisms—individual recognition, reciprocal credence, and cost–return sensitivity—as simulateable behavioral primitives for modeling the emergence of economic exchange;
@@
   • Reinterpret key findings from behavioral economics—such as inequity aversion, reciprocal cooperation, and altruistic punishment—as consistent expressions of underlying biological mechanisms, rather than anomalies to rational choice theory.
+Unlike prior models that treat cooperation heuristically or assume trust as exogenous, our framework offers a cognitively minimal and simulateable account of how such expectations emerge from social interaction.
 This work reframes economic modeling not as a top-down implementation of rules, but as a bottom-up emergence of institutions from cognitive interaction.
@@
 Our framework builds on these findings by situating primate reciprocity within a cognitive-behavioral model.
-We argue that capacities such as individual recognition, memory of interaction history, and fairness sensitivity form a biologically grounded substrate upon which more complex human economic systems could emerge.
+We argue that capacities such as individual recognition, memory of interaction history, and cost-return sensitivity form a behaviorally grounded substrate upon which more complex human economic systems could emerge.
 2.3 Social Exchange Without Markets
@@
 In contrast, our framework treats these preferences as emergent outcomes of specific cognitive-social mechanisms: individual recognition, memory of interaction, and cost–return evaluation.
-By shifting the focus from outcomes to underlying capacities, we aim to bridge behavioral economic observations with a biologically grounded account of exchange behavior.
+By shifting the focus from outcomes to underlying capacities, we aim to bridge behavioral economic observations with a behaviorally grounded account of exchange behavior.
 Figure 2: Evidence from anthropology, primatology, and developmental psychology suggests that the behavioral foundations of economic exchange lie in reciprocity—not barter or market logic.
@@
 What emerges from these systems is not merely institutional scaffolding, but a structured behavioral logic—one that we formalize in the following section.
-3.2 The Three Core Mechanisms of Biological Reciprocity
+3.2 The Three Behavioral Mechanisms Enabling Reciprocity
 If the earliest forms of economic interaction were not driven by impersonal barter, then what behavioral dynamics made sustained exchange possible?
@@
   • Cost–return sensitivity: allowing agents to regulate cooperative behavior based on dynamic payoff asymmetries.
-These three capacities are supported by converging evidence from nonhuman primates and human infants, suggesting they are both biologically plausible and developmentally early.
-Importantly, they are not merely descriptive—they are simulateable as modular components in multi-agent systems.
-While primate studies have richly documented reciprocal behaviors—grooming, food sharing, coalition support—the cognitive prerequisites for sustaining such behavior are often left implicit.
+These mechanisms are supported by converging evidence from nonhuman primates and human infants, suggesting they are biologically plausible, developmentally early, and cognitively minimal.
+Each is also behaviorally tractable and simulateable in multi-agent systems—whether as memory modules, value estimators, or policy regulators.
+While primate studies have richly documented behaviors like grooming, food sharing, and coalition support, the cognitive prerequisites for sustaining these interactions are often left implicit.
 As (Schino & Aureli, 2010) note, reciprocity is frequently treated as a local response pattern, rather than a structured process involving memory and expectation.
@@
 Without recognition or anticipation of future interaction, trust-like behavior cannot functionally emerge.
-By defining these mechanisms as behavioral primitives, we offer a bottom-up model of economic emergence—one that grounds institutional systems not in rules, but in biologically plausible, cognitively minimal substrates of social interaction.
+By defining these three mechanisms as simulateable behavioral primitives, we provide a bottom-up framework for modeling the emergence of economic exchange—one that avoids presupposing institutions, markets, or symbolic trust, and instead builds from capacities that are both empirically grounded and computationally implementable.
+Figure 3: The agent integrates three input pathways via behaviorally grounded mechanisms—recognition, credence, and cost–return evaluation—to support dynamic decisions in social exchange contexts.
 4 Theoretical Framework: Simulateable Primitives for Economic Exchange
-Existing efforts in agent-based cooperation often rely on engineered incentives, fixed strategy spaces, or symbolic approximations of concepts like “trust.” What remains missing is a biologically grounded, simulateable account of the behavioral capacities that make reciprocal exchange possible.
-We propose three minimal mechanisms—individual recognition, reciprocal credence, and cost–return sensitivity—as a compositional framework for simulating the emergence of economic exchange.
-These mechanisms are:
-  • Cognitively minimal: each is supported by empirical data from nonhuman primates and human infants;
-  • Behaviorally tractable: each has observable correlates in real-world cooperation;
-  • Simulateable: each can be implemented in agent-based systems as a decision module or update function.
-This framework allows social exchange to be modeled from the bottom up—without relying on predefined markets, institutions, or payoff matrices.
 4.1 Individual Recognition
-We define individual recognition as the ability to identify and distinguish specific agents across time, enabling the tracking of prior social interactions.
-This capacity underlies the first and most basic source of reciprocal credence: memory of positive experiences with known others.
-Without recognition, no information about past cooperation can be meaningfully retained or used to guide future behavior.
+Definition 4.1 (Individual Recognition).
+Individual recognition is the capacity to identify and distinguish specific social partners across time.
+It enables agents to track prior interactions with particular individuals, forming the basis for memory-based expectations about future behavior.
+This capacity underlies the first and most basic source of reciprocal credence: without individual recognition, no information about past cooperation can be meaningfully retained or applied to guide future interaction.
+Empirical Support.
 Among nonhuman primates, this capacity is especially robust.
@@
 This conceptual ambiguity makes trust difficult to operationalize, especially in comparative, developmental, or artificial settings.
-To address this gap, we introduce the concept of reciprocal credence: a cognitively grounded, graded expectation that another agent will respond to cooperation with cooperation.
-Reciprocal credence does not assume moral virtue or social obligation.
-Instead, it arises from informational cues—typically based on one of two elements: (1) past prosocial behavior directed toward the self or others, or (2) contextual signals that make cooperative return likely.
-It is not a feeling, but a function: a decision-weighting variable that governs the willingness to enter or continue reciprocal exchange.
-Unlike categorical trust, reciprocal credence is inherently scalar: it reflects how likely an agent believes cooperation will be reciprocated, not whether it will be or not.
-Treating reciprocal credence as a scalar variable is essential to modeling cooperation as a dynamic, context-sensitive behavior.
-We distinguish three primary sources of reciprocal credence:
+To address this gap, we introduce the concept of reciprocal credence: a cognitively grounded, graded belief that another agent will respond to cooperation with cooperation.
+Unlike categorical trust, reciprocal credence is scalar, updateable, and context-sensitive.
+It is not a feeling, but a functional belief state—one that governs an agent’s willingness to initiate or sustain cooperative behavior.
+Definition 4.2 (Reciprocal Credence).
+Reciprocal credence is a scalar, updateable belief that another agent will respond to cooperation with cooperation.
+It is not a feeling, but a functional decision variable that governs an agent’s willingness to initiate or sustain cooperative behavior.
+It arises from two temporally grounded types of inference:
+  • Retrospective inference, based on memory of past prosocial behavior directed toward the self or others;
+  • Prospective inference, based on contextual signals that indicate future reciprocity is likely—such as repeated interaction, social roles, or institutional incentives.
+To make these sources operational in multi-agent systems, we distinguish three canonical informational pathways:
   1. 1.
-    Direct positive interaction history, where an individual has experienced prior cooperative behavior, prosocial engagement, or emotional bonding with another.
+    Direct interaction history, where an agent has experienced prior cooperative behavior, prosocial engagement, or affiliative signaling from a specific partner.
   2. 2.
-    Reputational inference, where the individual observes or learns about others’ cooperative behavior.
+    Reputational inference, based on observed or reported behavior toward others—supporting third-party inference of prosocial disposition.
   3. 3.
-    Role-based expectation, where social roles or institutional norms lead to default expectations of reciprocity.
-Each source varies in reliability, stability, and cognitive cost, but all contribute to the emergence of trust-like behavior without requiring moral commitment.
-Together, these sources provide the informational basis for reciprocal credence—a flexible, graded expectation that others will act in ways conducive to cooperation.
+    Role-based or contextual expectation, where social roles, environmental incentives, or normative structures lead to default assumptions that cooperation will be returned, even in the absence of personal history.
+These pathways differ in reliability, cognitive cost, and developmental trajectory—but together, they provide the informational grounding for reciprocal credence as a simulateable and biologically plausible substrate for reciprocal cooperation.
+Empirical Support.
 Empirical studies support the differential emergence of these sources across species and developmental stages.
@@
 A note on “trust”.
-It is also important to distinguish social trust from what might be better described as functional dependency.
-Contemporary systems such as ChatGPT or Google are often described as being “trusted,” when in reality they are simply relied upon due to their consistent output, broad adoption, and lack of failure.
-This perceived “trust” is better understood as a product of usability, network effects, and repeated exposure — not reciprocal social accountability.
-A similar semantic conflation is common in financial discourse, where “trust in money” or “trust in banks” is frequently cited as foundational to economic stability.
-However, what is usually meant in such contexts is not interpersonal trust, but confidence in the continuity of institutional enforcement, liquidity guarantees, or the absence of systemic failure.
+The term “trust” is often applied to systems like ChatGPT or Google, but what is described is not social trust—it is functional reliability through repeated exposure and low failure rates.
+These systems do not engage in reciprocal reasoning or uphold social obligations.
+This semantic conflation becomes especially problematic in high-stakes domains like finance and healthcare, where “trust” is better understood as a demand for exceptional systemic robustness—not social accountability.
+In finance, phrases like “trust in banks” or “trust in money” reflect belief in institutional continuity, enforcement mechanisms, and liquidity guarantees—not interpersonal confidence.
+What matters is structural resilience against systemic collapse.
+In healthcare, “trust in doctors” or “hospital systems” similarly refers to confidence in diagnostic accuracy, institutional safeguards, and low error rates.
+Patients rely on these systems not because of perceived benevolence, but because failures are rare and tightly bounded.
 4.3 Cost–return sensitivity
-We define cost–return sensitivity as the capacity to evaluate whether engaging in cooperative behavior yields a net benefit over time.
+Definition 4.3 (Cost–Return Sensitivity).
+Cost–return sensitivity is the capacity to evaluate whether engaging in cooperative behavior yields a net benefit over time.
 Unlike categorical heuristics (e.g., always cooperate or always defect), this sensitivity enables agents to condition their participation on the payoff structure of ongoing interactions.
-It allows withdrawal from exploitative relationships and reinforcement of beneficial ones, serving as a regulatory mechanism for reciprocal dynamics.
+It supports both the withdrawal from exploitative relationships and the reinforcement of beneficial ones, serving as a regulatory mechanism for reciprocal dynamics.
 This capacity need not involve explicit calculation; approximate assessments—such as tracking recent imbalances in giving and receiving—are sufficient to support adaptive behavior.
-For example, chimpanzees are less likely to groom or share with partners who have failed to reciprocate in the past, even if those costs are not immediate or symmetrical.
-Similarly, infants gradually reduce helping behavior toward agents who consistently fail to reciprocate attention or aid.
-Such sensitivity ensures that reciprocal exchange remains robust under uncertainty, preventing exploitation while maintaining flexibility in dynamic social environments.
+Empirical Support.
 In nonhuman primates, cost–return sensitivity manifests in both prosocial and retaliatory contexts.
@@
 As a building block of economic cognition, it governs the micro-dynamics of exchange long before the emergence of price or property.
-5 Possible Implementation via Memory Design
-While each proposed mechanism is tractable in isolation, integrating them into a unified architecture remains an open challenge—particularly due to current limitations in memory design for LLM-based agents.
-Most existing multi-agent frameworks lack persistent, structured memory, resulting in shallow social modeling and limited capacity for behavioral adaptation.
-We propose that each behavioral primitive—individual recognition, reciprocal credence, and cost–return sensitivity—can be realized through targeted extensions to agent memory and retrieval mechanisms:
-  • Individual recognition can be implemented using identity-specific memory slots or embedding-indexed retrieval systems, enabling agents to store and retrieve interaction histories with named partners.
-This supports continuity in agent behavior and the ability to condition future responses on partner identity.
-  • Reciprocal credence may be represented as an internal latent variable or scalar annotation associated with each agent in memory (e.g., “Agent X: credence = 0.7”), and updated through lightweight heuristics (e.g., recency-weighted prosocial acts).
-This allows LLM agents to infer, represent, and update partner expectations without requiring explicit reward functions.
-  • Cost–return sensitivity can be implemented through structured memory logs of past cooperative attempts and outcomes.
-Agents may track cumulative give–take asymmetries and adjust future investments accordingly.
-Prompts can surface this information as part of decision rationale (e.g., “Agent Y has taken without giving for three turns”).
-These strategies enable biologically inspired reciprocity to be modeled using memory-centric extensions, rather than reinforcement learning or symbolic planning.
-By shifting the burden of social behavior onto structured memory design and prompt-level reasoning, this framework opens a tractable path toward more socially competent LLM agents.
+5 Simulation Pathways and Implementation
+We define three simulateable behavioral primitives—individual recognition, reciprocal credence, and cost–return sensitivity—that can serve as the minimal cognitive capacities required for scalable reciprocal behavior in multi-agent systems.
+Each can be approximated through lightweight memory architectures and prompt-level reasoning, without relying on explicit reinforcement learning.
+  • Individual recognition: Agents should be able to differentiate between social partners and retrieve information specific to each.
+This may reflect accumulated positive history, social familiarity, or bonded interaction patterns—not necessarily explicit naming.
+  • Reciprocal credence: Agents should maintain an internal estimate of how likely a partner is to return cooperation.
+This estimate integrates: (1) prior prosocial behavior directed toward the agent or observed in others, and (2) contextual signals that make cooperative return likely.
+The value is asymmetric, updateable, and sensitive to both interaction and situation.
+  • Cost–return sensitivity: Agents should track asymmetries in past exchanges, allowing them to adjust future cooperative investment.
+This mechanism supports dynamic calibration of effort and helps stabilize long-term reciprocity.
+These primitives do not prescribe a specific architecture, but define necessary behavioral functions for any agent seeking to simulate socially grounded exchange.
+By shifting the modeling burden toward structured memory and belief inference, they offer a tractable and biologically plausible path for implementing reciprocal behavior in LLM-based systems.
 6 Implications and Discussion
@@
 Institutional Complexity.
-Although we argue that modern exchange institutions build upon a biologically grounded behavioral substrate, we do not claim that such mechanisms alone are sufficient to explain the emergence of formal systems like currency, interest-bearing debt, or taxation.
+Although we argue that modern exchange institutions build upon a behaviorally grounded behavioral substrate, we do not claim that such mechanisms alone are sufficient to explain the emergence of formal systems like currency, interest-bearing debt, or taxation.
 These likely require additional cultural, historical, and symbolic developments.
@@
 6.2 Theoretical Implications
+Reciprocity as the Foundation of Social Structure.
+The term cooperation is frequently used as a catch-all for prosocial behavior, but its vagueness makes it difficult to model or implement.
+Without a clear cognitive basis, many systems rely on artificial payoff matrices, fixed strategies, or exogenous incentives—mechanisms that simulate outcomes but not the behavioral processes behind them.
+In contrast, reciprocity offers a cognitively grounded and behaviorally structured alternative.
+It is not merely repeated cooperation, but a mechanism based on individual recognition, interaction history, and cost–return regulation.
+These features make reciprocity tractable, simulateable, and aligned with real-world behavior.
+More importantly, reciprocity underlies many core functions of modern society.
+From social favors and mentorships to business alliances and venture capital, cooperation often depends on asymmetric, memory-based, and deferred exchanges.
+Even scientific systems—such as peer review, citation, and referrals—operate through implicit reciprocity, not formal enforcement.
+We argue that reciprocity—not abstract cooperation—is the minimal behavioral substrate of social structure.
+It enables agents to build conditional expectations, regulate ongoing interactions, and scale bilateral exchange into stable institutional patterns.
+Far from being a subset of cooperation, reciprocity is the mechanism that makes cooperation sustainable and socially extensible in the first place.
 From Institutions to Cognition.
@@
 We argue that this vagueness limits its explanatory value.
-To address this, we introduce reciprocal credence: a graded, biologically grounded expectation that cooperation will be returned.
+To address this, we introduce reciprocal credence: a graded, behaviorally grounded expectation that cooperation will be returned.
 It emerges from recognition, interaction history, and cost–return sensitivity—not from moral obligation.
@@
 Memory and bonding may be embedded within the learning processes that govern credence formation.
-With the rise of large language model (LLM)-based agents and embodied multi-agent simulators (e.g., Habitat, ThreeDWorld, AutoGen), it is now possible to explore whether agents endowed with these primitives can spontaneously develop systems of reciprocal exchange, role specialization, symbolic debt, or proto-monetary behaviors.
+With the rise of large language model (LLM)-based agents and embodied multi-agent simulators (e.g., VirtualHome (Puig et al., 2018), ThreeDWorld (Gan et al., 2020), AutoGen (Wu et al., 2023)), it is now possible to explore whether agents endowed with these primitives can spontaneously develop systems of reciprocal exchange, role specialization, symbolic debt, or proto-monetary behaviors.
 Such experiments would help identify the minimal substrate for scalable cooperation and provide a generative framework for building socially competent artificial agents.
@@
 Future work can embed these primitives into large-scale multi-agent environments to explore the spontaneous formation of scalable cooperation and exchange systems.
+Together, these primitives offer not just theoretical insight, but a tractable interface for embedding socially competent behavior in next-generation agent systems.
 Impact Statement

### v3

@@
 Abstract
-A key challenge in multi-agent AI is modeling social cooperation under realistic behavioral constraints.
-Many foundational concepts in economics and ethics—such as ”trust” or ”morality”—are often defined informally, without operational criteria or cognitive grounding, which limits their testability and implementation in artificial agents.
-Drawing on converging empirical evidence from primate behavior, infant cognition, and economic anthropology, we propose a conceptual framework composed of three cognitively minimal mechanisms: individual recognition, reciprocal credence, and cost–return sensitivity.
-This framework reframes trust as a graded cognitive expectation, providing a simulateable basis for reciprocal exchange in artificial agents, enabling the bottom-up emergence of scalable cooperation, proto-economic norms, and institutional structure.
+The origins of economic behavior remain unresolved—not only in the social sciences but also in AI, where dominant theories often rely on predefined incentives or institutional assumptions.
+Contrary to the longstanding myth of barter as the foundation of exchange, converging evidence from early human societies suggests that reciprocity—not barter—was the foundational economic logic, enabling communities to sustain exchange and social cohesion long before formal markets emerged.
+Yet despite its centrality, reciprocity lacks a simulateable and cognitively grounded account.
+Here, we introduce a minimal behavioral framework based on three empirically supported cognitive primitives—individual recognition, reciprocal credence, and cost–return sensitivity—that enable agents to participate in and sustain reciprocal exchange, laying the foundation for scalable economic behavior.
+These mechanisms scaffold the emergence of cooperation, proto-economic exchange, and institutional structure from the bottom up.
+By bridging insights from primatology, developmental psychology, and economic anthropology, this framework offers a unified substrate for modeling trust, coordination, and economic behavior in both human and artificial systems.
 Machine Learning, ICML
@@
 1 Introduction
-A persistent gap in economic simulation is the lack of behavioral grounding: most models begin with institutions—markets, contracts, payoff matrices—while assuming the cognitive and social mechanisms that make such systems viable.
-This limits both the explanatory power and simulateability of artificial economic agents.
-This issue is mirrored in broader AI+Society research.
-Concepts like “trust,” “value,” and “cooperation” are often treated as abstract labels, operationalized only at the outcome level.
-Without grounding these constructs in observable mechanisms, models risk simulating the surface of economic life without modeling its origins.
-In this paper, we argue that the foundations of economic exchange—and by extension, scalable cooperation—can be traced to three cognitively minimal, behaviorally grounded mechanisms: individual recognition, reciprocal credence, and cost–return sensitivity.
-These capacities are observed across human and nonhuman primates, and support sustained reciprocal exchange without requiring symbolic trust or formal enforcement.
-Anthropological accounts have long shown that early human exchange was not rooted in barter or equivalence, but in socially embedded reciprocity.
-Systems like the Kula ring (Malinowski, 2013) emphasized obligation and alliance over transaction.
-Similarly, primatology has documented robust reciprocal behaviors among chimpanzees and bonobos—grooming (De Waal, 1997), food sharing, and coalition—indicating that the behavioral foundations of reciprocal exchange are observable across species and precede formal market systems.
-We propose a modular framework built from these three mechanisms, designed to serve as behavioral primitives in economic simulation.
-This approach not only bridges cognitive science and institutional theory, but also enables bottom-up modeling of how debt, credit, and monetary behavior may emerge from repeated social interaction.
+The standard origin story of economics begins with barter: individuals trading goods directly, with money and markets emerging to reduce friction.
+Yet this narrative is a myth.
+Ethnographic and historical research finds little evidence for barter as a primary mode of early exchange (Sahlins, 2013; Mauss, 2024; Malinowski, 2013).
+Instead, early human societies were structured around reciprocity—a temporally extended system of giving, receiving, and returning, embedded in social relationships.
+Surprisingly, most economic models still begin where barter leaves off: with institutions, contracts, and payoff matrices.
+These frameworks assume—but do not explain—the behavioral substrate that makes cooperation possible in the first place.
+As a result, we lack simulateable accounts of how economic behavior emerges from interaction among agents.
+This blind spot extends to broader AI + society research.
+Core concepts like “trust,” “value,” and “cooperation” are often used without precise definitions or behavioral grounding.
+Rather than being modeled as well-defined, simulateable mechanisms, they are treated as vague abstractions—obscuring rather than explaining the dynamics of social interaction.
+In this paper, we argue that the foundations of economic exchange—and by extension, scalable cooperation—emerge not from barter or institutional design, but from reciprocity: a structured pattern of interaction built on three cognitively minimal mechanisms— individual recognition, reciprocal credence, and cost–return sensitivity.
+These mechanisms, observed across humans and nonhuman primates, support partner-specific cooperation without relying on symbolic trust, contracts, or centralized enforcement.
+We formalize these mechanisms as simulateable behavioral primitives, providing a bottom-up framework grounded in cognitive and behavioral evidence.
+By integrating insights from anthropology, developmental psychology, and economic anthropology, our approach offers a biologically grounded alternative to institution-first models of cooperation.
 Our contribution.
-We offer a theoretical realignment: a cognitively grounded account of reciprocal exchange that clarifies how economic systems can emerge from social behavior rather than institutional design.
-We argue that reciprocity is not just a component of cooperation, but the minimal cognitive substrate from which stable social structures emerge.
+We provide a minimal theoretical framework grounded in behaviorally observable mechanisms. This framework yields four core contributions:
   • Identify three minimal mechanisms—individual recognition, reciprocal credence, and cost–return sensitivity—as simulateable behavioral primitives for modeling the emergence of economic exchange;
@@
   • Reframe “trust” as a scalar, simulateable expectation—reciprocal credence—rather than a moral abstraction;
-  • Reinterpret key findings from behavioral economics—such as inequity aversion, reciprocal cooperation, and altruistic punishment—as consistent expressions of underlying biological mechanisms, rather than anomalies to rational choice theory.
-Unlike prior models that treat cooperation heuristically or assume trust as exogenous, our framework offers a cognitively minimal and simulateable account of how such expectations emerge from social interaction.
-This work reframes economic modeling not as a top-down implementation of rules, but as a bottom-up emergence of institutions from cognitive interaction.
+  • Reinterpret classic behavioral findings—such as framing effects, loss aversion, and anchoring bias—as consistent expressions of bounded, biologically grounded cognition, rather than anomalies to rational choice theory.
+Unlike models that treat trust heuristically or assume cooperation as exogenous, our framework offers a minimal, simulateable foundation for how cooperative expectations and exchange structures emerge from interaction.
+This allows us to reframe the origins of economic systems not as top-down implementations of institutions, but as bottom-up constructions grounded in cognitive behavior.
 Ethical Statement
@@
 2.1 Agent-Based Social Simulation
-Recent advances in multi-agent language models have produced systems capable of planning, negotiation, and scripted cooperation (Park et al., 2023; Li et al., 2023).
-However, these models typically lack behavioral grounding: they simulate interaction as token-level dialogue without modeling the internal mechanisms that make social exchange stable—such as partner memory, reciprocity tracking, or cost–benefit sensitivity.
-While some frameworks include persistent memory modules or heuristic-based reasoning, these components often lack principled constraints.
-Agents either memorize indiscriminately or update state in ways that are unstable or opaque, making it difficult to simulate long-term reciprocal dynamics.
-As a result, apparent cooperation often emerges from prompt bias or hardcoded behavior, not from simulateable social inference.
-This gap has two key consequences: (1) It limits the explanatory value of multi-agent simulations as models of human social behavior, and (2) it prevents systematic exploration of how early exchange mechanisms scale into institutional forms.
-Our framework addresses this by identifying minimal, biologically plausible mechanisms that can be embedded in LLM agents to support grounded, dynamic reciprocity over time.
+Recent work on multi-agent language models has enabled scripted cooperation and task planning (Park et al., 2023; Li et al., 2023), but these systems rarely model the cognitive mechanisms that support stable social interaction—such as partner tracking, interaction memory, or sensitivity to prior outcomes.
+While some frameworks add memory or heuristics, these are typically unconstrained and ad hoc, limiting their ability to simulate reciprocal dynamics over time.
+Apparent cooperation often stems from prompt bias or hardcoded behavior rather than simulateable social inference.
+Our framework addresses this gap by identifying minimal, biologically grounded mechanisms that support dynamic reciprocity, enabling more realistic modeling of social behavior and scalable exchange.
 2.2 Reciprocal Behavior in Nonhuman Primates
-Research in primatology provides critical insight into the behavioral foundations of reciprocity.
-Species such as chimpanzees and bonobos engage in reciprocal behaviors across contexts like food sharing, grooming, and coalition support.
-For instance, de Waal (De Waal, 1997) illustrated how chimpanzees form long-term social alliances maintained through reciprocal exchange, and Brosnan et al.
-(Brosnan & De Waal, 2003) demonstrated that primates are sensitive to fairness and inequity in outcomes.
-While these studies have richly documented the behavioral forms of reciprocity, they often stop short of linking such behaviors to the broader architecture of economic interaction.
-The question of how these mechanisms scale or evolve into structured systems of exchange remains relatively underexplored.
-Our framework builds on these findings by situating primate reciprocity within a cognitive-behavioral model.
-We argue that capacities such as individual recognition, memory of interaction history, and cost-return sensitivity form a behaviorally grounded substrate upon which more complex human economic systems could emerge.
+Studies in primatology reveal that species such as chimpanzees and bonobos engage in reciprocal acts across grooming, food sharing, and alliance formation.
+For instance, de Waal (De Waal, 1997) documented long-term cooperative relationships maintained through delayed reciprocity, while Brosnan et al.
+(Brosnan & De Waal, 2003) showed sensitivity to fairness and outcome inequity.
+Although these findings offer strong evidence of reciprocal behavior, they are seldom linked to broader economic structures or coordination dynamics.
+The underlying cognitive mechanisms—and how they scale to support complex social systems—remain largely unexplored.
 2.3 Social Exchange Without Markets
-Anthropological accounts have long challenged the barter-origin myth that underpins classical economic theory.
-Foundational works such as The Gift (Mauss, 2024) and Stone Age Economics (Sahlins, 2013) argue that early exchange was not driven by market logic, but embedded in webs of social relationships, mutual obligation, and symbolic prestige.
-These accounts have been instrumental in shifting attention away from transactional equivalence and toward socially embedded forms of value.
-Yet, they often bracket the question of what kinds of cognitive and behavioral mechanisms make such systems feasible.
-In other words: what allows individuals to engage in sustained, deferred, and socially contingent exchange?
-Our framework complements these anthropological insights by identifying the minimal behavioral architecture that enables reciprocity to stabilize across time and social partners.
-We argue that reciprocity is not merely a cultural invention, but a behaviorally tractable structure observable across species.
-2.4 From Behavioral Anomalies to Cognitive Substrates
-Behavioral economics has significantly advanced our understanding of social preferences, particularly in contexts of fairness, cooperation, and reciprocity.
-Paradigms such as the Ultimatum Game (Güth et al., 1982), Trust Game (Berg et al., 1995), and Public Goods Game (Fehr & Gächter, 2002) reveal systematic deviations from self-interested rationality, with participants rejecting unfair offers, rewarding cooperation, and punishing free-riders.
-These empirical patterns have led to influential models of “inequity aversion” (Fehr & Schmidt, 1999) and “strong reciprocity” (Gintis, 2000), which posit internalized fairness norms and preferences for norm enforcement.
-However, these models typically treat the relevant capacities—such as fairness sensitivity or social tracking—as given, rather than as targets of explanation.
-In contrast, our framework treats these preferences as emergent outcomes of specific cognitive-social mechanisms: individual recognition, memory of interaction, and cost–return evaluation.
-By shifting the focus from outcomes to underlying capacities, we aim to bridge behavioral economic observations with a behaviorally grounded account of exchange behavior.
+Anthropological work has challenged the classical view that early economies began with barter.
+Foundational accounts, such as Mauss’s The Gift (Mauss, 2024) and Sahlins’s Stone Age Economics (Sahlins, 2013), emphasize that early exchange was embedded in networks of kinship, obligation, and prestige—rather than driven by market logic or equivalent trade.
+These studies highlight the importance of social context and long-term reciprocity, but often leave open the question of what mechanisms make such systems sustainable.
+Specifically, the cognitive conditions under which deferred, partner-contingent exchange can stabilize across time are rarely formalized.
 Figure 2: Evidence from anthropology, primatology, and developmental psychology suggests that the behavioral foundations of economic exchange lie in reciprocity—not barter or market logic.
@@
 3.1 The Myth of Barter and the Reciprocal Foundations of Exchange
-The textbook account of economic origins often begins with barter: the notion that early humans exchanged goods directly in the absence of currency, with markets and money arising later to reduce transactional friction.
-This narrative, while intuitively appealing, has little support in either ethnographic or historical records.
-As Sahlins argued in Stone Age Economics (Sahlins, 2013), early exchange systems were not structured around equivalence, but around relational modes of reciprocity—generalized, balanced, and negative.
-These forms differ not in rationality but in social distance and temporal orientation: generalized reciprocity, typical among kin, involved open-ended giving without expectation of immediate return; balanced reciprocity, among peers, entailed roughly equivalent exchanges over time; and negative reciprocity, often among rivals or strangers, reflected opportunistic or exploitative behavior.
-A paradigmatic ethnographic example is the Kula ring, described by Malinowski (Malinowski, 2013), in which ceremonial armbands and necklaces circulate across Trobriand Islands through long-term, prestige-based exchanges.
-These objects held little utilitarian value, yet were exchanged with strict directional patterns, social memory, and partner exclusivity.
-What appears as a symbolic ritual is, upon closer inspection, a complex system of relational tracking, reputational assessment, and obligation extension—precisely the kinds of cognitive demands that structure early forms of exchange.
-Mauss’s The Gift (Mauss, 2024) further formalized this insight by showing how gift-giving in archaic societies is embedded in a triadic obligation: to give, to receive, and to reciprocate.
-The gift is not merely a transfer of value, but a means of sustaining long-term social bonds.
-When viewed from this angle, early exchange is not the product of impersonal rational choice, but of biologically and socially grounded behaviors—tracking individuals, remembering interactions, maintaining alliances, and adjusting for asymmetrical returns.
-These foundations suggest that exchange predates markets not just historically, but cognitively.
-What emerges from these systems is not merely institutional scaffolding, but a structured behavioral logic—one that we formalize in the following section.
+The standard narrative of economic origins begins with barter: the idea that early humans exchanged goods directly, with markets and money emerging later to reduce friction.
+But this account lacks support in ethnographic and historical records.
+As Sahlins argued in Stone Age Economics (Sahlins, 2013), early exchange was not based on equivalence but on relational reciprocity: generalized reciprocity involved open-ended giving among kin; balanced reciprocity entailed delayed equivalence among peers; and negative reciprocity reflected opportunism among strangers.
+Malinowski’s Kula ring (Malinowski, 2013) exemplifies long-term, prestige-based exchange using non-utilitarian objects, governed by memory, exclusivity, and directional flow.
+Beneath the ritual lies a cognitively rich structure of tracking, obligation, and reputation.
+Mauss’s The Gift (Mauss, 2024) formalized this logic as a triadic obligation—to give, receive, and reciprocate—framing gift exchange as a cognitive and social mechanism for sustaining long-term bonds.
+Together, these accounts suggest that early human societies were not built on barter, but on reciprocity: a temporally extended structure of interaction grounded in memory, obligation, and repeated engagement.
+Far from a peripheral exchange strategy, reciprocity was the behavioral infrastructure of early economies—the cognitive and social foundation that enabled resource flow, social cohesion, and relational continuity.
+It sustained cooperation long before the emergence of contracts or currency, and predates markets not only historically, but cognitively—anchored in the minimal behavioral logic we formalize in the next section.
 3.2 The Three Behavioral Mechanisms Enabling Reciprocity
-If the earliest forms of economic interaction were not driven by impersonal barter, then what behavioral dynamics made sustained exchange possible?
-We argue that reciprocal exchange emerged not from a need to equalize value, but from the ability to maintain accountability and continuity across repeated social interactions.
-These early exchanges were not market transactions, but embedded social patterns shaped by memory, mutual responsiveness, and partner-specific interaction history.
-Exchange, in this view, arises not from price formation, but from behavioral infrastructure: the cognitive and social capacities that allow individuals to track others, assess cooperative value, and anticipate reciprocity.
-To formalize these behavioral foundations, we propose a cognitively grounded framework composed of three simulateable mechanisms:
-  • Individual recognition: enabling agents to identify and re-engage specific social partners over time;
-  • Reciprocal credence: a scalar, updateable expectation that cooperation will be returned;
-  • Cost–return sensitivity: allowing agents to regulate cooperative behavior based on dynamic payoff asymmetries.
-These mechanisms are supported by converging evidence from nonhuman primates and human infants, suggesting they are biologically plausible, developmentally early, and cognitively minimal.
-Each is also behaviorally tractable and simulateable in multi-agent systems—whether as memory modules, value estimators, or policy regulators.
-While primate studies have richly documented behaviors like grooming, food sharing, and coalition support, the cognitive prerequisites for sustaining these interactions are often left implicit.
-As (Schino & Aureli, 2010) note, reciprocity is frequently treated as a local response pattern, rather than a structured process involving memory and expectation.
-Yet reciprocal exchange, by definition, requires continuity.
-Without recognition or anticipation of future interaction, trust-like behavior cannot functionally emerge.
-By defining these three mechanisms as simulateable behavioral primitives, we provide a bottom-up framework for modeling the emergence of economic exchange—one that avoids presupposing institutions, markets, or symbolic trust, and instead builds from capacities that are both empirically grounded and computationally implementable.
-Figure 3: The agent integrates three input pathways via behaviorally grounded mechanisms—recognition, credence, and cost–return evaluation—to support dynamic decisions in social exchange contexts.
+If reciprocity—not barter—underpins early economic life, its roots must predate formal institutions.
+This raises a fundamental question: what minimal behavioral and cognitive mechanisms are sufficient to sustain reciprocal exchange?
+Nonhuman primates offer a critical comparative lens.
+In chimpanzees and bonobos, sustained, partner-contingent behaviors—such as food sharing, grooming, and coalition support—exhibit structured patterns of reciprocity and serve as core mechanisms for maintaining group cohesion and social stability (De Waal, 1997).
+Comparative psychology isolates the minimal cognitive substrates that support such behaviors.
+By examining species with simpler interaction structures, it avoids cultural and institutional confounds—revealing the baseline capacities sufficient for reciprocity to emerge and stabilize.
+To formalize these foundations, we introduce a cognitively grounded framework comprising three simulateable mechanisms:
+  • Individual recognition: identifying and re-engaging specific social partners over time;
+  • Reciprocal credence: an updateable expectation that cooperation will be returned;
+  • Cost–return sensitivity: modulating cooperative behavior in response to payoff asymmetries.
+We treat these as simulateable behavioral primitives—providing a bottom-up foundation for modeling the emergence of economic exchange, without presupposing institutions, markets, or symbolic trust.
+Figure 3: The agent integrates three input pathways via behaviorally grounded mechanisms—recognition, credence, and cost–return evaluation—to support dynamic decisions in reciprocal exchange.
 4 Theoretical Framework: Simulateable Primitives for Economic Exchange
@@
 Definition 4.1 (Individual Recognition).
-Individual recognition is the capacity to identify and distinguish specific social partners across time.
-It enables agents to track prior interactions with particular individuals, forming the basis for memory-based expectations about future behavior.
-This capacity underlies the first and most basic source of reciprocal credence: without individual recognition, no information about past cooperation can be meaningfully retained or applied to guide future interaction.
+Individual recognition is the capacity to identify specific social partners over time, enabling agents to track past interactions and form expectations about future behavior.
+It is the prerequisite for memory-based reciprocity: without it, no past cooperation can inform future decisions.
 Empirical Support.
-Among nonhuman primates, this capacity is especially robust.
-A recent study shows that chimpanzees and bonobos remember familiar individuals after decades of separation, and preferentially attend to those with whom they had positive social histories (Lewis et al., 2023).
-These findings show that individual recognition in apes is long-term and identity-specific, enabling the tracking of social partners across time.
-In nonhuman primates, individual recognition enables partner-specific cooperation.
-In rope-pulling tasks, chimpanzees adjusted their actions based on the identity and prior behavior of their partners—pausing, glancing, or gesturing to initiate coordination (Hirata & Fuwa, 2007).
-Dyads with higher food-sharing tolerance were also more successful in joint tasks, indicating that cooperative memory is structured around known social partners(Melis et al., 2006b).
-In humans, this ability emerges early.
-Infants as young as 14 months selectively help or share with individuals who have previously acted prosocially (Dunfield et al., 2011).
-Such findings suggest that recognition, even in its early form, is tied to social history and used to guide future cooperative behavior.
+Nonhuman primates show strong individual recognition. Chimpanzees and bonobos remember familiar individuals after decades and preferentially attend to former social partners (Lewis et al., 2023).
+In cooperative tasks, chimpanzees adjust behavior based on partner identity and prior interactions—initiating coordination through glances, pauses, or approach behaviors—and succeed more often with tolerant, familiar partners (Hirata & Fuwa, 2007; Melis et al., 2006b).
+In humans, this ability emerges in infancy.
+By 14 months, children selectively help those who previously acted prosocially (Dunfield et al., 2011), suggesting that recognition is tied to social memory and used to guide future cooperation.
 4.2 Reciprocal Credence
 The term “trust” is widely used across disciplines, yet rarely defined with precision.
-In both everyday language and scientific literature, it is invoked to describe moral commitment, emotional closeness, institutional reliability, and behavioral expectation—often interchangeably.
-This conceptual ambiguity makes trust difficult to operationalize, especially in comparative, developmental, or artificial settings.
-To address this gap, we introduce the concept of reciprocal credence: a cognitively grounded, graded belief that another agent will respond to cooperation with cooperation.
-Unlike categorical trust, reciprocal credence is scalar, updateable, and context-sensitive.
-It is not a feeling, but a functional belief state—one that governs an agent’s willingness to initiate or sustain cooperative behavior.
+It is invoked to describe moral commitment, emotional closeness, institutional reliability, and behavioral expectation—often interchangeably.
+This ambiguity makes it difficult to operationalize trust, especially in comparative, developmental, or artificial settings.
+To address this, we introduce the concept of reciprocal credence:
 Definition 4.2 (Reciprocal Credence).
-Reciprocal credence is a scalar, updateable belief that another agent will respond to cooperation with cooperation.
-It is not a feeling, but a functional decision variable that governs an agent’s willingness to initiate or sustain cooperative behavior.
+Reciprocal credence is a graded, updateable belief that cooperative behavior will be reciprocated.
+It functions as a decision variable governing an agent’s willingness to initiate or sustain cooperation.
 It arises from two temporally grounded types of inference:
   • Retrospective inference, based on memory of past prosocial behavior directed toward the self or others;
-  • Prospective inference, based on contextual signals that indicate future reciprocity is likely—such as repeated interaction, social roles, or institutional incentives.
+  • Prospective inference, based on contextual signals that indicate future reciprocity is likely—such as social roles or reputational cues.
 To make these sources operational in multi-agent systems, we distinguish three canonical informational pathways:
@@
   2. 2.
-    Reputational inference, based on observed or reported behavior toward others—supporting third-party inference of prosocial disposition.
+    Third-party inference, based on observed or reported behavior toward others—supporting third-party inference of prosocial disposition.
   3. 3.
     Role-based or contextual expectation, where social roles, environmental incentives, or normative structures lead to default assumptions that cooperation will be returned, even in the absence of personal history.
-These pathways differ in reliability, cognitive cost, and developmental trajectory—but together, they provide the informational grounding for reciprocal credence as a simulateable and biologically plausible substrate for reciprocal cooperation.
+Together, these sources form a minimal, biologically grounded substrate for modeling reciprocal behavior in both human and artificial agents.
 Empirical Support.
-Empirical studies support the differential emergence of these sources across species and developmental stages.
-In chimpanzees, reciprocal cooperation is consistently observed when individuals have direct histories of prosocial interaction, with little evidence for reputation-based inference or role-based expectation.
-Similarly, human infants display early sensitivity to agents who have helped them directly, yet show limited use of third-party observation or role generalization.
-These patterns suggest that direct positive interaction is the foundational layer of reciprocal credence, while the other two sources emerge developmentally and culturally.
-Across species, direct prosocial interaction serves as the most robust and foundational basis for reciprocal credence.
-In wild chimpanzees, grooming is exchanged not merely in-the-moment, but over extended time horizons.
-Longitudinal data reveal that grooming given and received is significantly correlated across dyads, even with delays spanning several days—indicating that chimpanzees maintain memory of past interactions and calibrate future behavior accordingly (Gomes et al., 2009).
-Experimental studies further show that chimpanzees prefer to collaborate with previous cooperative partners, suggesting they encode individual-specific social history and deploy it strategically in cooperative contexts (Melis et al., 2006a).
-Human infants exhibit a comparable early capacity.
-By three years of age, children are less likely to help individuals who have previously acted with harmful intent—even when the harm was directed at a third party, and even if no harm was successfully done.
-This selective avoidance indicates that infants track not only outcomes but also the moral valence of agents’ prior actions when deciding to engage prosocially (Vaish et al., 2010).
-Together, these findings support the primacy of direct positive interaction in generating reciprocal credence.
-Unlike inference or abstraction, this source relies on lived experience—anchoring cooperative expectations in observable history, not hypothetical roles or indirect cues.
+Direct positive interaction provides the most robust foundation for reciprocal credence.
+In chimpanzees, grooming is exchanged not only immediately but across time: dyads reciprocate over multi-day delays, suggesting memory-based calibration of cooperation (Gomes et al., 2009).
+They also preferentially collaborate with partners who have helped them before (Melis et al., 2006a), indicating an ability to track and evaluate individual-specific cooperative history.
+Similarly, by age three, human children avoid helping agents who previously harmed others, even if the harm was unsuccessful (Vaish et al., 2010)—suggesting early sensitivity to past intent in guiding prosocial choice.
+By contrast, third-party inference and role-based expectation are rarely observed in nonhuman primates and emerge only later in human development.
+Nonetheless, both are deeply embedded in human social life, forming the basis for reputation systems, institutional roles, and generalized cooperation beyond direct experience.
 A note on “trust”.
@@
 Definition 4.3 (Cost–Return Sensitivity).
-Cost–return sensitivity is the capacity to evaluate whether engaging in cooperative behavior yields a net benefit over time.
-Unlike categorical heuristics (e.g., always cooperate or always defect), this sensitivity enables agents to condition their participation on the payoff structure of ongoing interactions.
-It supports both the withdrawal from exploitative relationships and the reinforcement of beneficial ones, serving as a regulatory mechanism for reciprocal dynamics.
-This capacity need not involve explicit calculation; approximate assessments—such as tracking recent imbalances in giving and receiving—are sufficient to support adaptive behavior.
+Cost–return sensitivity is the capacity to regulate cooperative behavior based on perceived payoff asymmetries over time.
+Rather than relying on fixed heuristics, agents adjust participation by tracking whether interactions yield sustained net benefit—enabling the avoidance of exploitation and reinforcement of beneficial exchange.
 Empirical Support.
-In nonhuman primates, cost–return sensitivity manifests in both prosocial and retaliatory contexts.
-Chimpanzees engage in grooming and service-like exchanges where helping behavior depends on previous benefits received from others (De Waal, 1997).
-They are also capable of negative reciprocity: retaliating against individuals who have stolen from them, even when such retaliation yields no direct benefit—suggesting an expectation of equitable return and a sensitivity to intentional harm (Jensen et al., 2007).
-In humans, early traces of this mechanism appear in infancy.
-By 18 months, children selectively direct helping or sharing behavior toward agents who previously acted fairly or cooperatively (House et al., 2013).
-These decisions are not based on abstract principles of justice, but on perceived patterns of cost and return—who gave, who withheld, and under what circumstances.
-Taken together, these findings suggest that cost–return sensitivity enables organisms to calibrate cooperative investment, maintaining reciprocity in the absence of formal institutions or explicit contracts.
-As a building block of economic cognition, it governs the micro-dynamics of exchange long before the emergence of price or property.
+In nonhuman primates, cost–return sensitivity appears in both prosocial and punitive contexts.
+Chimpanzees modulate grooming and cooperative behavior based on prior benefits received (De Waal, 1997), and retaliate against theft even without immediate gain—suggesting expectations of equity and sensitivity to intentional harm (Jensen et al., 2007).
+Human infants show comparable intuitions.
+By 18 months, they preferentially help agents who previously acted fairly or cooperatively (House et al., 2013), calibrating prosocial behavior not through abstract norms, but through observed patterns of giving and withholding.
+Together, these findings suggest that cost–return sensitivity enables organisms to regulate cooperative investment in the absence of formal contracts—providing a cognitive substrate for reciprocity long before the emergence of markets.
 5 Simulation Pathways and Implementation
@@
 This mechanism supports dynamic calibration of effort and helps stabilize long-term reciprocity.
-These primitives do not prescribe a specific architecture, but define necessary behavioral functions for any agent seeking to simulate socially grounded exchange.
-By shifting the modeling burden toward structured memory and belief inference, they offer a tractable and biologically plausible path for implementing reciprocal behavior in LLM-based systems.
+These primitives specify functional requirements rather than architectural constraints, offering a practical path toward socially grounded reciprocity in multi-agent systems.
 6 Implications and Discussion
@@
 Comparative Methodology.
-Comparative psychology and developmental psychology use fundamentally different experimental paradigms.
-While primate studies often capture interactive, ecologically valid behaviors, human infant research typically relies on simplified, highly constrained designs.
-This asymmetry limits the granularity with which we can compare capacities across species.
+Primate and infant studies rely on distinct paradigms: the former captures ecologically embedded behavior; the latter often uses constrained, artificial settings.
+This methodological gap limits fine-grained cross-species comparison of cognitive substrates.
 Institutional Complexity.
-Although we argue that modern exchange institutions build upon a behaviorally grounded behavioral substrate, we do not claim that such mechanisms alone are sufficient to explain the emergence of formal systems like currency, interest-bearing debt, or taxation.
-These likely require additional cultural, historical, and symbolic developments.
+We do not claim that our framework alone accounts for complex institutions such as money, debt, or taxation.
+These might require further symbolic, cultural, and historical layers beyond minimal reciprocity mechanisms.
 Social Scale and Enforcement.
-Our framework focuses on the dyadic and small-group foundations of reciprocal exchange, where individuals recognize and respond to the behavior of specific partners.
-It does not yet account for mechanisms that support large-scale, anonymous cooperation—such as third-party punishment, norm internalization, or reputation management across broader social networks.
-These are critical features of modern institutions, and extending the framework to address how such systems emerge from or layer upon basic social cognition remains a key direction for future work.
+The current framework addresses partner-contingent reciprocity in small groups.
+It does not yet explain population-scale mechanisms—such as third-party punishment or distributed reputation—that enable anonymous cooperation.
+Bridging this gap remains a key direction for extending the model.
 6.2 Theoretical Implications
@@
 In contrast, reciprocity offers a cognitively grounded and behaviorally structured alternative.
-It is not merely repeated cooperation, but a mechanism based on individual recognition, interaction history, and cost–return regulation.
-These features make reciprocity tractable, simulateable, and aligned with real-world behavior.
+Rather than relying on externally defined payoffs, it emerges from simple agent-level mechanisms that support dynamic, partner-contingent decision-making.
+This makes reciprocity tractable, simulateable, and aligned with real-world behavior.
 More importantly, reciprocity underlies many core functions of modern society.
@@
 We argue that this vagueness limits its explanatory value.
-To address this, we introduce reciprocal credence: a graded, behaviorally grounded expectation that cooperation will be returned.
-It emerges from recognition, interaction history, and cost–return sensitivity—not from moral obligation.
-In contrast, large-scale “trust” in systems like money or platforms is better understood as confidence in robustness, not social accountability.
-This reframing clarifies how reciprocal behavior scales without assuming institutional trust.
+To address this, we introduce reciprocal credence as a graded, behaviorally grounded expectation—specific to reciprocal exchange—that governs whether an agent initiates or maintains cooperation.
+It depends not on social bonding, but on observable interaction history and contextual inference.
+In contrast, large-scale “trust” in systems like money or platforms is better understood as confidence in system robustness—not interpersonal accountability or relational bonding.
+This reframing clarifies how large-scale exchange systems can emerge without relying on social-bond-based trust.
 Cross-Species Foundations of Exchange.
@@
 Unifying Behavioral Economics through Cognitive Foundations.
-Many findings in behavioral economics—such as fairness preferences, reciprocal cooperation, and sensitivity to framing or loss—are often treated as deviations from rational choice theory.
-Yet evidence from primatology and developmental psychology suggests that these behaviors reflect consistent biological patterns: they emerge early in human development, recur across great apes, and follow predictable social-cognitive constraints.
-Rather than anomalies or patchwork heuristics, these tendencies may be better understood as structured expressions of an underlying behavioral logic—one shaped by the cognitive demands of navigating social exchange.
+If economic behavior is an extension of biologically grounded reciprocity, then so-called “biases” identified in behavioral economics—such as anchoring bias, loss aversion are not anomalies—they are direct extensions of biological traits.
 6.3 Future Work
-6.3.1 Operationalizing Economic Emergence in Agent Simulations
-A key direction for future research is to instantiate the three proposed behavioral primitives—individual recognition, reciprocal credence, and cost–return sensitivity—as explicit, modular components in LLM-based multi-agent systems.
-This involves not only defining each mechanism’s internal representation (e.g., persistent memory slots for identity recognition, scalar credence values for reciprocity tracking, and interaction-weighted cost–return summaries), but also studying how they interact dynamically across time and partners.
-We propose that future simulations systematically test:
-  • How memory constraints (e.g., limited identity tracking) affect the emergence of sustained cooperation;
-  • Whether agents with scalar credence variables spontaneously differentiate partners and regulate prosocial behavior;
-  • Under what conditions cost–return asymmetries trigger withdrawal, retaliation, or symbolic substitutes (e.g., tokens, credit markers).
-Such experiments can help identify the minimal memory and inference structure necessary to support scalable exchange.
-More broadly, this line of work offers a concrete path toward simulating institutional emergence—from micro-level reciprocity to meso-level norms and proto-economic artifacts—grounded in biologically plausible cognition.
-6.3.2 Simulating Institution Formation via Behavioral Primitives
-A promising direction for future work is to simulate the emergence of economic institutions—such as debt, symbolic exchange, and trust networks—by embedding behavioral primitives into artificial agents.
-These primitives, derived from comparative psychology and developmental research, include individual recognition, reciprocal credence, and cost–return sensitivity—as the minimal substrate for modeling scalable cooperation.
-Memory and bonding may be embedded within the learning processes that govern credence formation.
-With the rise of large language model (LLM)-based agents and embodied multi-agent simulators (e.g., VirtualHome (Puig et al., 2018), ThreeDWorld (Gan et al., 2020), AutoGen (Wu et al., 2023)), it is now possible to explore whether agents endowed with these primitives can spontaneously develop systems of reciprocal exchange, role specialization, symbolic debt, or proto-monetary behaviors.
-Such experiments would help identify the minimal substrate for scalable cooperation and provide a generative framework for building socially competent artificial agents.
-Importantly, this approach also enables us to test under what conditions behavioral reciprocity transitions into institutionalization: for example, whether limited memory capacity leads to symbolic tokens, or whether increasing agent density necessitates generalized trust representations.
-Simulation environments offer a new way to bridge cognitive models of exchange with the design of future artificial social systems.
-6.3.3 Exchange among unfamiliar agents
-One unresolved but critical question concerns the possibility of reciprocal exchange among unfamiliar agents.
-While most existing studies—whether in primatology, developmental psychology, or economic anthropology—focus on cooperation within socially familiar contexts, little is known about how exchange unfolds in the absence of recognition, shared norms, or interaction history.
-Transactions between unfamiliar individuals, especially without institutional enforcement, may require distinct cognitive or communicative scaffolding that is not yet well understood.
-We highlight this not as a limitation of the current framework, but as an open frontier that future research must address.
+Operationalizing Economic Emergence.
+A critical next step is to embed our proposed primitives—individual recognition, reciprocal credence, and cost–return sensitivity—as explicit modules in multi-agent simulations.
+Future experiments should systematically test how memory constraints, scalar reciprocity estimates, and payoff asymmetries influence the emergence of sustained cooperation and reciprocal exchange structures.
+Institution Formation via Behavioral Mechanisms.
+Further simulations should explore whether agents equipped with these minimal cognitive primitives spontaneously develop proto-institutions such as symbolic debt, token systems, or role-based cooperation.
+Advances in embodied simulators (e.g., VirtualHome (Puig et al., 2018), AutoGen (Wu et al., 2023)) offer opportunities to bridge cognitive theories of exchange with practical implementations of scalable social structures.
+Reciprocity Among Unfamiliar Agents.
+Finally, investigating reciprocal exchange among unfamiliar agents—without prior recognition or shared interaction history—represents an important frontier.
+Understanding the cognitive scaffolds enabling such interactions could significantly advance our grasp of decentralized cooperation mechanisms.
 7 Conclusion
-This paper proposes a cognitively grounded framework for modeling the behavioral origins of economic exchange in multi-agent systems.
-Instead of relying on institutional assumptions or symbolic abstractions, we identify three simulateable primitives—individual recognition, reciprocal credence, and cost–return sensitivity—that support sustained reciprocal cooperation in the absence of formal mechanisms.
-Our framework addresses a central modeling challenge: how to operationalize constructs such as “trust” or “morality,” which are foundational in human societies but often lack consensus definitions or concrete implementation paths in computational systems.
-Drawing on evidence from primate behavior, infant cognition, and economic anthropology, we reframe these abstract concepts in terms of observable, cognitively minimal mechanisms—providing a tractable basis for implementation in artificial agents.
-This perspective encourages a shift from top-down economic modeling to bottom-up simulation, where complex structures emerge from agent interaction rather than being imposed.
-Future work can embed these primitives into large-scale multi-agent environments to explore the spontaneous formation of scalable cooperation and exchange systems.
-Together, these primitives offer not just theoretical insight, but a tractable interface for embedding socially competent behavior in next-generation agent systems.
+Contrary to the classical narrative of barter, converging evidence from early human societies suggests that economic life began with reciprocity—long-term, partner-contingent exchange embedded in social relationships.
+This behavioral substrate enabled communities to circulate resources, maintain cohesion, and sustain cooperation well before markets or institutions emerged.
+Yet despite its foundational role in early economies, reciprocity remains under-formalized in both economic theory and computational modeling.
+We propose a cognitively grounded framework that identifies three simulateable mechanisms—individual recognition, reciprocal credence, and cost–return sensitivity—as the minimal substrate for scalable exchange.
+By grounding abstract terms like “trust” and “cooperation” in observable mechanisms, this approach shifts economic modeling from top-down design to bottom-up simulation.
+These primitives not only clarify the origins of exchange, but also provide a tractable substrate for implementing reciprocal behavior in multi-agent systems.
 Impact Statement
-This paper proposes a cognitively grounded framework for simulating the behavioral origins of economic exchange, with potential applications in multi-agent modeling, artificial social systems, and foundational research in AI + Society.
-While the work is theoretical and does not involve data collection or deployment, it provides biologically plausible primitives that could improve the interpretability and social competence of agent-based systems.
-We foresee no foreseeable negative societal consequences or ethical risks associated with this research.
+This paper introduces a cognitively grounded framework for simulating the behavioral foundations of economic exchange, based on biologically plausible primitives.
+By shifting the focus from top-down incentives to bottom-up social cognition, this approach enables more realistic and interpretable models of reciprocal cooperation.
+We anticipate that this framework could significantly improve agent-based economic simulations, particularly in multi-agent systems concerned with trust, fairness, and coordination.
+While the work is theoretical and does not involve deployment or data collection, we are not aware of any negative societal or ethical risks associated with this research.